A projectile is thrown upward so that its distance above the ground after t seconds is h = - 16t2 + 440t.

After how many seconds does it reach its maximum height?


help me please
1
Expert's answer
2014-06-05T04:37:14-0400
The trajectory of its motion is a parabola with downward branches (the coefficient with t^2 less then 0)
The maximum height is in the vertex of this parabola. Using the formula t0 = -b/2a we get
tmax = -440/(-2*16) = 13.75 seconds
Answer: 13.75 seconds.